# Solving the Cold Start Problem on One’s Own as an End User via Preference Transfer

### v1

@@ 6.3 Case Study
Table 2: Case Study. The source items the user interacted and the $K=20$ target items selected by Pretender. Pretender selects the very items that the user interacted with in the source service as shown in red. The user interacted with movies in 1996 and 1997, and Pretender mostly selects movies in 1996 and 1997. Notable exceptions are Alice in Wonderland (1951) and Fantasia (1940) selected by Pretender, as shown in blue. We hypothesis that this reflects the fact that the user thumbed down Cats Don’t Dance (1997), which is an animated movie for children. Since Cats Don’t Dance (1997) is not in the source service, Pretender selects similar movies from the target service and sends a signal that the user does not like this type of movies.
@@ 7 Conclusion
    • We showed that the regret converges at a rate of $O(\frac{1}{\sqrt{K}})$ for MMD (Corollary 3.6) and $O(K^{\frac{1}{d+2}})$ for the Wasserstein distance (Theorem 3.15).
@@ Acknowledgements
We thank Yuki Takezawa for helpful discussions.
@@
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{j,j^{\prime}}\mathbb{E}_{I}\left[\frac{I_{j}}{K}\frac{I_{j^{\prime}}}{K}\right]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-2\sum_{j,j^{\prime}}\mathbb{E}_{I}\left[\frac{I_{j}}{K}\hat{{\bm{w}}}_{j^{\prime}}\right]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle+\sum_{j,j^{\prime}}\mathbb{E}_{I}[\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (46)
=$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\frac{1}{K^{2}}\sum_{j,j^{\prime}}\mathbb{E}_{I}[I_{j}I_{j^{\prime}}]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-2\sum_{j,j^{\prime}}\frac{1}{K}\mathbb{E}_{I}[I_{j}]\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle+\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (47)
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\frac{1}{K^{2}}\sum_{j,j^{\prime}}\mathbb{E}_{I}[I_{j}I_{j^{\prime}}]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-2\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle+\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (48)
@@
=$\stackrel{{\scriptstyle\text{(d)}}}{{=}}\frac{1}{K^{2}}\sum_{j}\mathbb{E}_{I}[I_{j}]\langle\phi(x_{j}),\phi(x_{j})\rangle+\frac{1}{K^{2}}\sum_{j\neq j^{\prime}}\mathbb{E}_{I}[I_{j}]\mathbb{E}_{I}[I_{j^{\prime}}]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (51)
=$\stackrel{{\scriptstyle\text{(e)}}}{{=}}\frac{1}{K^{2}}\sum_{j}\mathbb{E}_{I}[I_{j}]\langle\phi(x_{j}),\phi(x_{j})\rangle+\sum_{j\neq j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (52)
@@
=$\stackrel{{\scriptstyle\text{(f)}}}{{=}}\frac{1}{K}\sum_{j}\hat{{\bm{w}}}_{j}\langle\phi(x_{j}),\phi(x_{j})\rangle-\sum_{j}\hat{{\bm{w}}}_{j}^{2}\langle\phi(x_{j}),\phi(x_{j})\rangle$ (54)
@@
=$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\frac{B}{K},$ (58)
@@
Let $\tilde{{\bm{w}}}^{\prime}_{j}=\frac{1}{K}$ if the $j$-th item is in the final output $\mathcal{D}_{T}$ and 00 otherwise. We have
@@
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\frac{\sqrt{B}}{K}\left|K-\sum_{j=1}^{2m}I_{j}\right|,$ (67)
@@
[…]
@@
Let $\mathcal{F}_{\text{Lip}}$ be the set of $1$-Lipschitz functions with $f(\frac{\mathbb{1}}{2})=0$. $\mathcal{F}_{\text{Lip}}$ can be $L_{\infty}$-covered by ~$\tilde{\mathcal{F}}_{\varepsilon}$, i.e., for any $f\in\mathcal{F}_{\text{Lip}}$, there exists $\tilde{f}\in\tilde{\mathcal{F}}_{\varepsilon}$ such that $\|f-\tilde{f}\|_{\infty}\leq\varepsilon$, with $|f(x)|\leq\frac{\sqrt{d}}{2},\forall f\in\tilde{\mathcal{F}}_{\varepsilon}$ and |~$|\tilde{\mathcal{F}}_{\varepsilon}|\leq\exp\left(\log(3)\left(\frac{2\sqrt{d}}{\varepsilon}\right)^{d}\right)$.
@@
We devide the domain $[0,1]^{d}$ into $\lceil\frac{\sqrt{d}}{\varepsilon}\rceil^{d}$ hypercubes with diameter at most $\varepsilon$. For each hypercube $\mathcal{C}$ and $f\in\mathcal{F}_{\text{Lip}}$, let $f_{\mathcal{C}}=\frac{\max_{x\in\mathcal{C}}f(x)+\min_{x\in\mathcal{C}}f(x)}{2}$ and $\tilde{f}_{\mathcal{C}}$ be its nearest point in $\{k\varepsilon\mid k\in\mathbb{Z}\}$. We then define $\tilde{f}(x)=\tilde{f}_{\mathcal{C}}$ for $x\in\mathcal{C}$. We have $|f(x)-\tilde{f}(x)|\leq\varepsilon$ since the diameter of $\mathcal{C}$ is at most $\varepsilon$ and $f$ is $1$-Lipschitz. We count the number of possible $\tilde{f}$. For the hypercube $\mathcal{C}_{\text{center}}$ with $\frac{\mathbb{1}}{2}\in\mathcal{C}_{\text{center}}$, $\tilde{f}_{\mathcal{C}_{\text{center}}}=0$ since $f(\frac{\mathbb{1}}{2})=0$. Given the value of $\tilde{f}_{\mathcal{C}}$, possible values of an adjacent hypercube $\mathcal{C}^{\prime}$ are $\tilde{f}_{\mathcal{C}}-\varepsilon,\tilde{f}_{\mathcal{C}},\tilde{f}_{\mathcal{C}}+\varepsilon$. We can reach all the hypercubes by repeating this process. Therefore, the possible values of $\tilde{f}$ are at most $3^{s-1}$, where $s=\lceil\frac{\sqrt{d}}{\varepsilon}\rceil^{d}$ is the number of hypercubes. When $\frac{\sqrt{d}}{\varepsilon}\geq 1$, we have $\lceil\frac{\sqrt{d}}{\varepsilon}\rceil\leq\frac{\sqrt{d}}{\varepsilon}+1\leq\frac{2\sqrt{d}}{\varepsilon}$, and |~$|\tilde{\mathcal{F}}|\leq\exp\left(\log(3)\left(\frac{2\sqrt{d}}{\varepsilon}\right)^{d}\right)$. When $\frac{\sqrt{d}}{\varepsilon}<1$, we have |~$|\tilde{\mathcal{F}}|=1$, in which case the lemma holds trivially. ∎
@@
Let $\tilde{{\bm{w}}}^{\prime}_{j}=\frac{1}{K}$ if the $j$-th item is in the final output $\mathcal{D}_{T}$ and 00 otherwise. For any $\varepsilon>0$,
@@
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\frac{d}{4K},$ (94)
@@
By Lemma 3.12, |~$|\tilde{\mathcal{F}}_{\varepsilon}|\leq\exp\left(\log(3)\left(\frac{2\sqrt{d}}{\varepsilon}\right)^{d}\right)$. We have
@@
where (a) follows from Eq. 103. We set $\delta\leftarrow\delta/|\tilde{\mathcal{F}}|$ in Proposition 3.14 and apply the union bound, then, we have

### v2

@@ 6.3 Case Study
Table 2: Case Study. The source items the user interacted with and the $K=20$ target items selected by Pretender. Pretender selects the very items that the user interacted with in the source service as shown in red. The user interacted with movies in 1996 and 1997, and Pretender mostly selects movies in 1996 and 1997. Notable exceptions are Alice in Wonderland (1951) and Fantasia (1940) selected by Pretender, as shown in blue. We hypothesize that this reflects the fact that the user thumbed down Cats Don’t Dance (1997), which is an animated movie for children. Since Cats Don’t Dance (1997) is not in the target service, Pretender selects similar movies from the target service and sends a signal that the user does not like this type of movies.
@@ 7 Conclusion
    • We showed that the regret converges at a rate of $O(\frac{1}{\sqrt{K}})$ for MMD (Corollary 3.6) and $O(K^{-\frac{1}{d+2}})$ for the Wasserstein distance (Theorem 3.15).
@@ Acknowledgements
We thank Yuki Takezawa and Satoshi Hayakawa for helpful discussions.
@@
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{j,j^{\prime}}\mathbb{E}_{I}\left[\frac{I_{j}}{K}\frac{I_{j^{\prime}}}{K}\right]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-2\sum_{j,j^{\prime}}\mathbb{E}_{I}\left[\frac{I_{j}}{K}\hat{{\bm{w}}}_{j^{\prime}}\right]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle+\sum_{j,j^{\prime}}\mathbb{E}_{I}[\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (46)
$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\frac{1}{K^{2}}\sum_{j,j^{\prime}}\mathbb{E}_{I}[I_{j}I_{j^{\prime}}]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-2\sum_{j,j^{\prime}}\frac{1}{K}\mathbb{E}_{I}[I_{j}]\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle+\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (47)
$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\frac{1}{K^{2}}\sum_{j,j^{\prime}}\mathbb{E}_{I}[I_{j}I_{j^{\prime}}]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-2\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle+\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (48)
@@
$\stackrel{{\scriptstyle\text{(d)}}}{{=}}\frac{1}{K^{2}}\sum_{j}\mathbb{E}_{I}[I_{j}]\langle\phi(x_{j}),\phi(x_{j})\rangle+\frac{1}{K^{2}}\sum_{j\neq j^{\prime}}\mathbb{E}_{I}[I_{j}]\mathbb{E}_{I}[I_{j^{\prime}}]\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (51)
$\stackrel{{\scriptstyle\text{(e)}}}{{=}}\frac{1}{K^{2}}\sum_{j}\mathbb{E}_{I}[I_{j}]\langle\phi(x_{j}),\phi(x_{j})\rangle+\sum_{j\neq j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle-\sum_{j,j^{\prime}}\hat{{\bm{w}}}_{j}\hat{{\bm{w}}}_{j^{\prime}}\langle\phi(x_{j}),\phi(x_{j^{\prime}})\rangle$ (52)
@@
$\stackrel{{\scriptstyle\text{(f)}}}{{=}}\frac{1}{K}\sum_{j}\hat{{\bm{w}}}_{j}\langle\phi(x_{j}),\phi(x_{j})\rangle-\sum_{j}\hat{{\bm{w}}}_{j}^{2}\langle\phi(x_{j}),\phi(x_{j})\rangle$ (54)
@@
$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\frac{B}{K},$ (58)
@@
Let $\tilde{{\bm{w}}}^{\prime}_{j}=\frac{1}{K}$ if the $j$-th item is in the final output $\mathcal{D}_{T}$ and 0 otherwise. We have
@@
$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\frac{\sqrt{B}}{K}\left|K-\sum_{j=1}^{2m}I_{j}\right|,$ (67)
@@
$=\left\|\sum_{j=1}^{2m}\tilde{{\bm{w}}}^{\prime}_{j}\phi(x_{j})-\int\phi(x)\,d\mu_{S}(x)\right\|_{\mathcal{H}}$ (71)
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\left\|\sum_{j=1}^{2m}\tilde{{\bm{w}}}^{\prime}_{j}\phi(x_{j})-\sum_{j=1}^{2m}\tilde{{\bm{w}}}_{j}\phi(x_{j})\right\|_{\mathcal{H}}+\left\|\sum_{j=1}^{2m}\tilde{{\bm{w}}}_{j}\phi(x_{j})-\sum_{j=1}^{2m}\hat{{\bm{w}}}_{j}\phi(x_{j})\right\|_{\mathcal{H}}+\left\|\sum_{j=1}^{2m}\hat{{\bm{w}}}_{j}\phi(x_{j})-\int\phi(x)\,d\mu_{S}(x)\right\|_{\mathcal{H}}$ (72)
@@
Let $\mathcal{F}_{\text{Lip}}$ be the set of $1$-Lipschitz functions with $f(\frac{\mathbb{1}}{2})=0$. $\mathcal{F}_{\text{Lip}}$ can be $L_{\infty}$-covered by $\tilde{\mathcal{F}}_{\varepsilon}$, i.e., for any $f\in\mathcal{F}_{\text{Lip}}$, there exists $\tilde{f}\in\tilde{\mathcal{F}}_{\varepsilon}$ such that $\|f-\tilde{f}\|_{\infty}\leq\varepsilon$, with $|f(x)|\leq\frac{\sqrt{d}}{2},\forall f\in\tilde{\mathcal{F}}_{\varepsilon}$ and $|\tilde{\mathcal{F}}_{\varepsilon}|\leq\exp\left(\log(3)\left(\frac{2\sqrt{d}}{\varepsilon}\right)^{d}\right)$.
@@
We devide the domain $0,1]^{d}$ into $\lceil\frac{\sqrt{d}}{\varepsilon}\rceil^{d}$ hypercubes with diameter at most $\varepsilon$. For each hypercube $\mathcal{C}$ and $f\in\mathcal{F}_{\text{Lip}}$, let $f_{\mathcal{C}}=\frac{\max_{x\in\mathcal{C}}f(x)+\min_{x\in\mathcal{C}}f(x)}{2}$ and $\tilde{f}_{\mathcal{C}}$ be its nearest point in $\{k\varepsilon\mid k\in\mathbb{Z}\}$. We then define $\tilde{f}(x)=\tilde{f}_{\mathcal{C}}$ for $x\in\mathcal{C}$. We have $|f(x)-\tilde{f}(x)|\leq\varepsilon$ since the diameter of $\mathcal{C}$ is at most $\varepsilon$ and $f$ is $1$-Lipschitz. We count the number of possible $\tilde{f}$. For the hypercube $\mathcal{C}_{\text{center}}$ with $\frac{\mathbb{1}}{2}\in\mathcal{C}_{\text{center}}$, $\tilde{f}_{\mathcal{C}_{\text{center}}}=0$ since $f(\frac{\mathbb{1}}{2})=0$. Given the value of $\tilde{f}_{\mathcal{C}}$, possible values of an adjacent hypercube $\mathcal{C}^{\prime}$ are $\tilde{f}_{\mathcal{C}}-\varepsilon,\tilde{f}_{\mathcal{C}},\tilde{f}_{\mathcal{C}}+\varepsilon$. We can reach all the hypercubes by repeating this process. Therefore, the possible values of $\tilde{f}$ are at most $3^{s-1}$, where $s=\lceil\frac{\sqrt{d}}{\varepsilon}\rceil^{d}$ is the number of hypercubes. When $\frac{\sqrt{d}}{\varepsilon}\geq 1$, we have $\lceil\frac{\sqrt{d}}{\varepsilon}\rceil\leq\frac{\sqrt{d}}{\varepsilon}+1\leq\frac{2\sqrt{d}}{\varepsilon}$, and $|\tilde{\mathcal{F}}|\leq\exp\left(\log(3)\left(\frac{2\sqrt{d}}{\varepsilon}\right)^{d}\right)$. When $\frac{\sqrt{d}}{\varepsilon}<1$, we have $|\tilde{\mathcal{F}}|=1$, in which case the lemma holds trivially. ∎
@@
Let $\tilde{{\bm{w}}}^{\prime}_{j}=\frac{1}{K}$ if the $j$-th item is in the final output $\mathcal{D}_{T}$ and 0 otherwise. For any $\varepsilon>0$,
@@
$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\frac{d}{4K},$ (94)
@@
By Lemma 3.12, $|\tilde{\mathcal{F}}_{\varepsilon}|\leq\exp\left(\log(3)\left(\frac{2\sqrt{d}}{\varepsilon}\right)^{d}\right)$. We have
@@
[…]